Integration of [(a√(a-x))/√(a+x)]
1
Expert's answer
2020-05-04T18:02:06-0400

"I=\int{\frac{a\sqrt{a-x}}{\sqrt{a+x}}}dx=a\int{\sqrt{\frac{a-x}{a+x}}}dx"

Let

"\frac{a-x}{a+x}=t^2\\
a-x=t^2(a+x)\\
x=a\cdot\frac{1-t^2}{1+t^2}\\
dx=\frac{-4at}{(t^2+1)^2}dt"

"I=-4a^2\int{\frac{t^2}{(t^2+1)^2}}dt"

Let

"t=\tan y\\
dt=\frac{dy}{\cos^2y}\\
I=-4a^2\int{\frac{\tan^2y}{(\tan^2y+1)^2}\frac{dy}{\cos^2y}}=\\
=-4a^2\int{\sin^2ydy}=-2a^2\int{(1-\cos2y)dy}=\\
=-2a^2(y-\frac{\sin2y}{4})+C"


"y=\arctan t, t=\sqrt{\frac{a-x}{a+x}}\\
y=\arctan\sqrt{\frac{a-x}{a+x}}"

then

"I=-2a^2(\arctan\sqrt{\frac{a-x}{a+x}}-\frac{\sin(2\arctan\sqrt{\frac{a-x}{a+x}})}{4})+C"